# No Black Boxes: Interpretable and Interactable Predictive Healthcare with Knowledge-Enhanced Agentic Causal Discovery

### v1

@@ C Pseudocodes
3:$q_{search}=LLM_{search}(\mathcal{D}_{p},\mathcal{S}_{p},\mathcal{K}_{\Gamma})$
@@
10:Output: Summarized document set $\Gamma_{p}^{summary}=a_{\gamma}\mid\gamma\in\Gamma^{q}$.
Algorithm 2 Causal Discovery Agent, $\mathcal{A}_{causal}$.
Input: Empty graph $\mathcal{G}^{\emptyset}$ consists of entities from $\mathcal{D}_{p}\cup\mathcal{S}_{p}$, Candidate diseases , Summarized document set $\Gamma_{p}^{summary}$, diagnosis matrix $\mathbf{A}_{D}$
2:$t\triangleq 0$
@@
4:$\mathcal{G}^{s}_{t=0}=\text{LLM}_{\text{hypo-gen}}(\mathcal{G}^{\emptyset},\Gamma_{p}^{\text{summary}})$
@@
Algorithm 3 Decision-Making Agent, $\mathcal{A}_{decision}$.
@@
3:$\mathcal{D}_{\text{pred}},\mathcal{E}=\text{LLM}_{\text{decison}}(\mathcal{S}_{p},\Gamma_{p}^{\text{summary}},\mathcal{G}^{s},\mathcal{C})$
@@
1:Require: Pretrained LLM model, EHR training data $\mathcal{D}_{p}\mid p\in\mathcal{P}_{train}$, Knowledge vector database $\Gamma$ with meta data $\mathcal{K}_{\Gamma}$.
2:// Data processing
3:Calculate $\mathbf{A}_{T}$ and $\mathbf{A}_{D}$ with Equation 1 and 3.
4:// Inference
5:For $p\in\mathcal{P}_{test}$:
6:    Input: Diagnosis history $\mathcal{D}_{p}$
7:    Obtain candidate disease $\mathcal{S}_{p}$ with Equation 4.
8:    // Knowledge Synthesis Agent
9:    $\Gamma_{p}^{summary}=\mathcal{A}_{summary}(\mathcal{D}_{p},\mathcal{S}_{p},\mathcal{K}_{\Gamma})$
10:    // Causal Discovery Agent
11:    $\mathcal{G}^{s}=\mathcal{A}_{causal}(\mathcal{D}_{p},\mathcal{S}_{p},\Gamma_{p}^{summary},\mathbf{A}_{D})$
12:    // Decision-Making Agent
13:    $\mathcal{D}_{pred},\mathcal{E}=\mathcal{A}_{decision}(\mathcal{D}_{p}\mathcal{S}_{p},\Gamma_{p}^{summary},\mathcal{G}^{s},\mathcal{C})$
14:   Output: Final causal graph $\mathcal{G}^{s}$, predicted
15:   diagnosis $\mathcal{D}_{\text{pred}}$ and explanations $\mathcal{E}$.
D Prompt Details

### v2

@@ B Pseudocodes
[…]
@@
10:Output: Summarized document set $\Gamma_{p}^{\text{summary}}=a_{\gamma}\mid\gamma\in\Gamma^{q}$.
Algorithm 2 Causal Discovery Agent, $\mathcal{A}_{\text{causal}}$.
Input: Empty graph $\mathcal{G}^{\emptyset}$ consists of entities from $\mathcal{D}_{p}\cup\mathcal{S}_{p}$, Candidate diseases , Summarized document set $\Gamma_{p}^{\text{summary}}$, diagnosis matrix $\mathbf{A}_{D}$
[…]
@@
[…]
@@
Algorithm 3 Decision-Making Agent, $\mathcal{A}_{\text{dm}}$.
@@
[…]
@@
Require: Pretrained LLM model, EHR training data $\mathcal{D}_{p}\mid p\in\mathcal{P}_{tr}$, Knowledge vector database $\Gamma$ with meta data $\mathcal{K}_{\Gamma}$.
// Data processing
Calculate $\mathbf{A}_{T}$ and $\mathbf{A}_{D}$ with Eqs. 1 and 3.
// Inference
For $p\in\mathcal{P}_{test}$:
Input: Diagnosis history $\mathcal{D}_{p}$
  Obtain candidate disease $\mathcal{S}_{p}$ with Eq. 4.
// Knowledge Synthesis Agent
$\Gamma_{p}^{\text{summary}}=\mathcal{A}_{\text{knowledge}}(\mathcal{D}_{p},\mathcal{S}_{p},\mathcal{K}_{\Gamma})$
// Causal Discovery Agent
$\mathcal{G}^{s}=\mathcal{A}_{\text{causal}}(\mathcal{D}_{p},\mathcal{S}_{p},\Gamma_{p}^{\text{summary}},\mathbf{A}_{D})$
// Decision-Making Agent
$\mathcal{D}_{\text{pred}},\mathcal{E}=\mathcal{A}_{\text{dm}}(\mathcal{D}_{p}\mathcal{S}_{p},\Gamma_{p}^{\text{summary}},\mathcal{G}^{s},\mathcal{C})$
Output: Final causal graph $\mathcal{G}^{s}$, predicted
 diagnosis $\mathcal{D}_{\text{pred}}$ and explanations $\mathcal{E}$.
C Prompt Details